OpenStudy (anonymous):

Will someone please help me?!? Justify if completing the square is a good method for solving when the Discriminant is negative. Use any of your three functions as an example and respond in complete sentences.

OpenStudy (anonymous):

The function I created I want to use is --> \[f(x)=(x-3)(x-2)\]

OpenStudy (anonymous):

@mathmale @austinL @maxdrizner

OpenStudy (anonymous):

@phi can you please help I don't understand

OpenStudy (maxdrizner):

i honestly dont know how to do this,,, sorry

OpenStudy (anonymous):

Thank you anyway.

OpenStudy (phi):

The question says *** Justify if completing the square is a good method for solving when the Discriminant is negative. *** the discriminant is b^2 - 4ac where you a,b and c are the coefficients of the quadratic, when it is written in standard form: \[ y = a x^2 + bx + c \]

OpenStudy (anonymous):

I already know the standard form of my first function is \[y=x ^{2}-5x+6\]

OpenStudy (anonymous):

Kid STOP! I am trying to learn something.

OpenStudy (anonymous):

I blocked him, sorry

OpenStudy (phi):

what is the discriminant of your function ?

OpenStudy (anonymous):

I'm not sure what the discriminant is.

OpenStudy (phi):

the discriminant is b^2 - 4*a*c in your case a=1 (number in front of x^2) b = -5 and c= 6

OpenStudy (anonymous):

okay I'm following you.

OpenStudy (phi):

you would get 1 for the discriminant. I don't see how that helps us, because they are asking about a negative discriminant. for example, find the discriminant of \[ f(x) = x^2 -2x + 2\]

OpenStudy (phi):

can you find the discriminate of \[ f(x) = x^2 -2x + 2 \] a=1, b= -2 and c=2

OpenStudy (phi):

if you using completing the square on \[ x^2 -2x + 2 = 0 \\ x^2 -2x = -2 \\ x^2 -2x +(-2/2)^2 = -2 +(-2/2)^2 \\ x^2 -2x + 1 = -1 \\ (x-1)^2 = -1 \] now take the square root of both sides \[ \sqrt{(x-1)^2} = \sqrt{-1} \\ x-1 = \sqrt{-1} \\ x= 1 ± \sqrt{-1} \] That is the correct answer, but if you don't know about complex numbers, then you don't know how to take the square root of a negative number.

OpenStudy (anonymous):

I know \[\sqrt{-1} = i\]

OpenStudy (anonymous):

@phi

OpenStudy (phi):

So that was an example of using "complete the square" to solve a quadratic whose discriminant is negative. If none of your examples has a negative discriminant maybe you should add that example to your list.

OpenStudy (anonymous):

okay thank you. I really appreciate you helping me.

OpenStudy (anonymous):

can you help me with one more thing?? I need help finding the solutions of that same function. \[g(x)=(x-\sqrt{3})(x+\sqrt{3})\]

OpenStudy (anonymous):

\[x=\sqrt{3},-\sqrt{3}\]

OpenStudy (anonymous):

\[g(\sqrt{3})=(x-\sqrt{3})(x+\sqrt{3})\]

OpenStudy (anonymous):

\[g=\frac{ (x-\sqrt{3})(x+\sqrt{3}) }{ \sqrt{3} }\]

OpenStudy (phi):

When you find the solutions of a function, you need an equation. In other words if you have a function g(x) you could ask, "what x makes g(x) = 0 ? " or "what x makes g(x) = 3 ?" , etc. When you solve for the "roots of an equation" that means solve g(x) = 0

OpenStudy (phi):

*** I need help finding the solutions of that same function. \[ g(x)=(x-\sqrt{3})(x+\sqrt{3}) \] *** that is not an equation, that is a Definition... it explains that g(x) is short-hand for the stuff on the right side of the = sign. if you had, for example \[ (x-\sqrt{3})(x+\sqrt{3}) = 0 \] then the solutions would be x= -sqrt(3) and x= + sqrt(3)

OpenStudy (phi):

notice that finding the solutions is very easy, because f(x) is in factored form. If you started with \[ g(x) = x^2 -3 \] and had to solve for \[ x^2 -3 = 0 \] you normally would have to factor the left-hand side. But in this case, it is still easy to solve: \[ x^2 = 3 \\ x = ± \sqrt{3} \]

OpenStudy (phi):

btw, \( g(\sqrt{3} )\) is short-hand for replace "x" with \( sqrt{3} \) in the definition of g(x) you would say: \[ g(x)=(x-\sqrt{3})(x+\sqrt{3}) \\ g(\sqrt{3} )=(\sqrt{3} -\sqrt{3})(\sqrt{3} +\sqrt{3}) \\ g(\sqrt{3} )= 0 \cdot 2 \sqrt{3}\\ g(\sqrt{3} )= 0 \] you would also get 0 if you used - sqrt(3)